# Reachable Polyhedral Marching (RPM): An Exact Analysis Tool for Deep-Learned Control Systems

### v3

@@ I-A Overview
It is well known that ReLU networks implement continuous piecewise-affine (PWA) functions, that is, the input space for a ReLU network can be tessellated into polyhedra, and over each polyhedron the neural network is affine. The RPM algorithms explicitly finds this equivalent PWA representation for a given ReLU network. Figure 1 illustrates how RPM incrementally solves for this representation. The algorithm incrementally enumerates the polyhedra and affine pieces of the PWA function by solving a series of Linear Programs (LPs), and following analytical edge flipping rules to determine neighboring polyhedra. The algorithm starts with an initial polyhedron, then solves for its neighboring polyhedra, then neighbors of neighbors, etc., until the desired input set is tessellated. In this way, our method is geometrically similar to fast marching methods in optimal control [2], path planning [3, 4], and graphics [5, 6].
We then use computational methods for PWA reachability to compute forward and backward reachable sets for each polyhedron and associated affine function. The ability to compute forward and backward reachable sets allows us to search for and construct control invariant sets and ROAs without Lyapunov tools. Finally, we also propose an accelerated backward reachability procedure that leverages the connected enumeration of the RPM algorithm to restrict the number of polyhedra the algorithm searches over. We show that this accelerated procedure is guaranteed to enumerate the entire backward reachable set in the case that the deep network is a homeomorphism (a continuous bijection with continuous inverse). Checking this condition is simply done, and represents a novel general procedure for examining the invertibility of neural networks whose hidden layers are not bijections themselves.
@@ I-B Existing Work
The nnenum tool uses an exact reachability algorithm that is also incremental [10]. The difference between this tool and our RPM algorithm is that RPM enumerates affine regions of the neural network in a connected walk; each new region that is revealed is connected to a previous region. In contrast, the algorithm used by nnenum does not have this property; whether a new region is connected to a previous one is unknown by the algorithm. In Section VI we show how the connected enumeration property leads to an accelerated algorithm for computing ROAs. Another way to conceptualize this difference between RPM and nnenum is that RPM can also return the graph that describes which affine regions neighbor one another.
@@ I-C Contributions and Organization
In this paper, we build off the RPM algorithm to construct control invariant sets and ROAs for dynamical systems represented as neural networks. By taking this approach, we are able to compute intricate invariant sets that may be non-convex and even disconnected. In numerical examples, we compute ROAs for a learned van der Pol oscillator, and show this computation enjoys a $15$x speedup because the learned dynamics are homeomorphic. We then pair RPM with the MPT3 [11] Matlab toolbox to find an a control invariant set for a learned torque-controlled pendulum. Finally, we apply our algorithm to find a set of states stabilized by an image-based airplane runway taxiing system, TaxiNet [12]. The closed-loop system is PWA with ${\sim}250,000$ regions, orders of magnitude larger than PWA systems for which other nonlinear stability analysis approaches have been demonstrated.
The contributions of this paper are
  • For ReLU neural networks, a theorem to analytically determine an input space polyhedron from its neighbor by flipping neuron activations in the network.
@@
  • An accelerated exact backward reachability algorithm for ReLU networks that leverages a novel method for checking the invertibility of the network.
This paper builds upon an earlier conference version [13], that first introduced the RPM algorithm to convert a ReLU network into a PWA representation. This paper improves upon [13] by (i) adding a formal proof of correctness for the main RPM algorithm, (ii) extending the RPM framework to compute control invariant sets and ROAs, and (iii) introducing an accelerated method for computing ROAs.
@@ II Related Work
Our RPM algorithm is an exact method, meaning that there is no conservatism in the reachable sets computed by our algorithm. Our algorithm inherits all the advantages of exact methods, but is unique in that it enumerates the reachable set in an incremental and connected walk. This makes our high-level procedure more similar to explicit model predictive control methods [49] than to other neural network reachability methods. Finally, all intermediate polyhedra and affine map matrices of layer-by-layer methods must be stored in memory until the algorithm terminates, whereas with incremental methods, once a polyhedron-affine map pair is computed it can be sent to external memory and only a binary vector (the neuron activations) needs to be stored to continue the algorithm. This is especially useful because no method has been shown to accurately estimate the number of regions without explicit enumeration. Unlike other incremental approaches, ours explores affine regions in a connected walk. As a point of distinction, this allows our algorithm to not only output the PWA representation, but for each affine region also output the neighboring affine regions.
In this paper we demonstrate how RPM can be used not only for finite-time reachability, but also for computing control invariant sets and ROAs. Computing these sets tends to be much harder than computing finite-time reachable sets, but the associated guarantees hold for all time, unlike with finite-time reachable sets. Few methods exist for computing invariant sets or ROAs for dynamical systems represented as neural networks, and most that do rely on an optimization-based search of a Lyapunov function. In [50], the authors learn a quadratic Lyapunov function using semidefinite programming that results in an ellipsoidal ROA, a drawback of which is that ellipsoids are not very expressive sets. In contrast, nonconvex ROAs can be computed by [51, 52, 53]. These methods learn Lyapunov functions for a given autonomous dynamical system and verify the Lyapunov conditions either using many samples and Lipschitz continuity or mixed integer programming. In [54], an optimization and sampling-based approach for synthesizing control invariant sets is given, but as with searching for Lyapunov functions, it may fail in finding a control invariant set when one exists. All of these methods inherit the drawbacks of Lyapunov-style approaches wherein a valid Lyapunov function may not be found when one exists, and certifying the Lyapunov conditions is challenging. In contrast to these methods, we explicitly solve for the PWA form of the dynamics, then find invariant sets using reachability methods rather than Lyapunov theory. A reachability-based approach like ours can be more constructive and reliable than Lyapunov approaches, as we show in our experiments. Lastly, a non-Lyapunov method is given in [55], where the focus is on continuous-time dynamical systems, whereas our focus is on discrete-time.
@@
In contrast to the Lyapunov approaches, the reachability approach requires the user to supply an initial ‘seed’ set of states that is a ROA and uses backward reachability to grow the size of the ROA. Finding a seed ROA can be difficult for general nonlinear systems, but for PWA systems can be easily found due to the dynamics being locally linear. Although reachability-based computation of invariant sets has been explored for PWA dynamical systems before [60], a novelty in our work is pairing these methods with our RPM algorithm so that they can apply to dynamical systems represented as neural networks. Our proposed method for finding ROAs of neural networks follows the backward reachability approach from the literature. In addition, we show that our backward reachability algorithm may be sped up considerably when restricting RPM to enumerate a connected backward reachable set (a feature enabled by the connected walk of the algorithm).
@@ VI-A Control Invariant Sets
In this subsection we discuss a strategy for computing control invariant sets once we have the PWA representation of a dynamical system $\mathbf{x}_{t+1}=F(\mathbf{x}_{t},\mathbf{u}_{t})$. This strategy for computing control invariant sets is implemented in MPT3 specifically for PWA dynamical systems [11]. First, we note that for some state transition functions $F$, there may not exist a control invariant set. Existence of such a set depends on the state domain $\mathcal{X}$ and the set of admissible control inputs $\mathcal{U}$, as well as the specific effect of control inputs on the state transitions given by $F$. If a control invariant set exists, it can be computed via the following recursion which uses backward reachability,
@@ VI-B Regions of Attraction
In this subsection we specifically consider the problem of computing invariant ROAs given a ReLU network that represents an autonomous dynamical system $\mathbf{x}_{t+1}=F(\mathbf{x}_{t})$. One of the most common goals in designing a control policy is to stabilize the system around some equilibrium. By identifying a ROA, we are identifying a set of states which is stabilized.
@@ VI-C An Accelerated Procedure
The backward reachability algorithm as previously presented can be made more efficient if RPM is restricted to only enumerated a connected backward reachable set. For instance, since we know the fixed point $\mathbf{x}_{eq}$ must be in the ROA, initializing RPM at the fixed point and then only computing a connected backward reachable set of the seed ROA can restrict the number of polyhedra explored by RPM significantly. In this case, we can alter the backward reachability algorithm to only add neighbor polyhedra to the working set if they are neighbors of a polyhedron known to be in the backward reachable set.
@@
Specific conditions for a PWA function to be homeomorphic are given in [71].
@@ VII Examples
We then find a seed ROA for the stable affine system as described in Section VI-B2. Next, we verify the network is a homeomorphism by evaluating the conditions in Section VI-C; from this we know that a connected set in the output space will have a connected preimage. The details of the learned network and computation time for these steps is summarized in Table I.
@@
Number of Affine Regions $248$ Regions
Time to Enumerate Affine Regions $2.45$ Seconds
Time to Check if Homeomorphism $0.04$ Seconds
Time to Find Fixed Point $0.28$ Seconds
Time to Find Seed ROA $1.08$ Seconds
@@
Steps # in ROA # Explored Time (s) # Explored Time (s)
5 $500$ $594$ $5$ $7362$ $98$
10 $2525$ $2735$ $42$ $33,946$ $627$
15 $6569$ $6898$ $157$ $85,217$ $2280$
20 $12,553$ $12,995$ $413$- --
25 $20,949$ $21,554$ $954$- --
30 $32,753$ $33,500$ $1956$- --
@@ VII-B Pendulum Example
as a discrete-time model where $m=1$kg, $g=9.81$m/s2, $l=1$m, and $\theta$ is measured clockwise from the upright position. Details of the learned dynamics network are in Table III.
@@
Number of Affine Regions $217$ Regions
Time to Enumerate Affine Regions $3$ Seconds
@@ VII-C Verifying Image-Based Control
In this example we apply the tools of Section VI to a larger dynamical system that models the closed-loop behavior of a taxiing airplane controlled from camera images. Our goal is to investigate the set of states that are stabilized by a learned image-based controller by modeling the closed-loop dynamical system with a single ReLU network, and applying the accelerated procedure from Section VI.
The specific system we consider was introduced in [12] where a taxiing airplane is regulated to the centerline of a runway via rudder-angle control input based on runway images taken from a camera on the wing. Because of the image observation, we cannot model the observation from physical principles. Instead, we modify the learned models in [12] to obtain a network that approximates the observed image given a state
@@
Next, we use Algorithm 4 to retrieve the explicit PWA function for $F_{cl}$ over the domain $p\in[-5,5]$ m., $\theta\in[-15^{\circ},15^{\circ}]$. The resulting PWA function has 244,920 regions and took 3 hours to compute. The PWA function was found to not be a homeomorphism. Computation time for these steps is summarized in Table V. Since the function is not homeomorphic, when applying the accelerated algorithm from Section VI-C, we are not guaranteed that the resulting set will be invariant, however, we are still guaranteed that all states within this set asymptotically converge to the stable fixed point.
@@ VIII Conclusions
We proposed the Reachable Polyhedral Marching (RPM) algorithm to incrementally construct an exact PWA representation of a ReLU network. RPM computes an explicit PWA function representation for a given ReLU network that can then be used to quickly find forward and backward reachable sets of ReLU networks. In addition, we demonstrated how, for nonlinear dynamical systems modeled as ReLU networks, the RPM tool can be leveraged to (i) compute intricate control invariant sets and (ii) certify stability and compute ROAs. RPM enumerates affine regions and reachable sets in an incremental and connected fashion, which we demonstrate can lead to massive acceleration when computing ROAs. Lastly, RPM also provides a novel way to determine invertibility of ReLU networks.

### v4

@@ I-A Overview
It is well known that ReLU networks implement continuous piecewise-affine (PWA) functions, that is, the input space for a ReLU network can be tessellated into polyhedra, and over each polyhedron the neural network is affine. The RPM algorithm explicitly finds this equivalent PWA representation for a given ReLU network. Figure 1 illustrates how RPM incrementally solves for this representation. The algorithm incrementally enumerates the polyhedra and affine pieces of the PWA function by solving a series of Linear Programs (LPs), and following analytical edge flipping rules to determine neighboring polyhedra. The algorithm starts with an initial polyhedron, then solves for its neighboring polyhedra, then neighbors of neighbors, etc., until the desired input set is tessellated. In this way, our method is geometrically similar to fast marching methods in optimal control [2], path planning [3, 4], and graphics [5, 6].
We then use methods for PWA reachability to compute forward and backward reachable sets for each polyhedron and associated affine function. Computing forward and backward reachable sets allows us to search for and construct control invariant sets and ROAs without Lyapunov tools. Finally, we also propose an accelerated backward reachability procedure that leverages the connected enumeration of the RPM algorithm to restrict the number of polyhedra the algorithm searches over. We show that this accelerated procedure is guaranteed to enumerate the entire backward reachable set in the case that the deep network is a homeomorphism (a continuous bijection with continuous inverse). Checking this condition is simple, and represents a novel general procedure for examining the invertibility of neural networks whose hidden layers are not invertible themselves.
@@ I-B Existing Work
The authors of [10] also propose an incremental approach to visiting each affine region in a ReLU network. The difference between this tool and our RPM algorithm is that RPM enumerates affine regions of the neural network in a connected walk; each new region that is revealed is connected to a previous region. In contrast, the algorithm used in [10] does not have this property; whether a new region is connected to a previous one is unknown by the algorithm. In Section VI we show how the connected enumeration property leads to an accelerated algorithm for computing ROAs. Another way to conceptualize this difference between RPM and [10] is that RPM can also return the graph that describes which affine regions neighbor one another. This feature is enabled by RPM’s unique property of removing redundant constraints in the input-space polyhedra, resulting in the minimal description of each affine region.
@@ I-C Contributions and Organization
In [11], the RPM algorithm was introduced and demonstrated on multi-step reachability tasks. In contrast, the contributions of this paper are
  • A formal proof of correctness for the RPM algorithm via a theorem to analytically determine an affine region from its neighbor by flipping neuron activations in ReLU networks.
@@
  • An accelerated exact backward reachability algorithm for ReLU networks that does not require enumerating all affine regions, leveraging a novel approach for checking the invertibility of the network.
Through these contributions, we are able to compute intricate invariant sets that may be non-convex and even disconnected. In numerical examples, we compute ROAs for a learned van der Pol oscillator, and show this computation enjoys a $15$x speedup because the learned dynamics are homeomorphic. We then pair RPM with the MPT3 [12] Matlab toolbox to find an a control invariant set for a learned torque-controlled pendulum. Finally, we apply our algorithm to find a set of states stabilized by an image-based airplane runway taxiing system, TaxiNet [13]. The closed-loop system is PWA with ${\sim}250,000$ regions, orders of magnitude larger than PWA systems for which other nonlinear stability analysis approaches have been demonstrated.
@@ II Related Work
Our RPM algorithm is an exact method, meaning that there is no conservatism in the reachable sets computed by our algorithm. Our algorithm inherits all the advantages of exact methods, but is unique in that it enumerates the reachable set in an incremental and connected walk. This makes our high-level procedure similar to explicit model predictive control methods [49]. Finally, all intermediate polyhedra and affine map matrices of layer-by-layer methods must be stored in memory until the algorithm terminates, whereas with incremental methods, once a polyhedron-affine map pair is computed it can be sent to external memory and only a binary vector (the neuron activations) needs to be stored to continue the algorithm. This is especially useful because no method has been shown to accurately estimate the number of regions without explicit enumeration. Unlike other incremental approaches, ours explores affine regions in a connected walk. As a point of distinction, this allows our algorithm to not only output the PWA representation, but for each affine region also output the neighboring affine regions.
In this paper we demonstrate how RPM can be used not only for finite-time reachability, but also for computing control invariant sets and ROAs. Computing these sets tends to be much harder than computing finite-time reachable sets, but the associated guarantees hold for all time. Few methods exist for computing invariant sets or ROAs for dynamical systems represented as neural networks, and most that do rely on an optimization-based search of a Lyapunov function. In [50], the authors learn a quadratic Lyapunov function using semidefinite programming that results in an ellipsoidal ROA, a drawback of which is that ellipsoids are not very expressive sets. In contrast, nonconvex ROAs can be computed by [51, 52, 53]. These methods learn Lyapunov functions for a given dynamical system and verify the Lyapunov conditions either using samples and Lipschitz continuity or mixed integer programming. In [54], an optimization and sampling-based approach for synthesizing control invariant sets is given, but as with searching for Lyapunov functions, it may fail in finding a control invariant set when one exists. All of these methods inherit the drawbacks of Lyapunov-style approaches wherein a valid Lyapunov function may not be found when one exists, and certifying the Lyapunov conditions is challenging. In contrast to these methods, we explicitly solve for the PWA form of the dynamics, then find invariant sets using reachability methods. A reachability-based approach like ours can be more constructive and reliable than Lyapunov approaches, as we show in our experiments. Lastly, a non-Lyapunov method is given in [55], where the focus is on continuous-time dynamical systems, whereas our focus is on discrete-time.
@@
In contrast to Lyapunov approaches, the reachability approach requires an initial ‘seed’ set of states that is a ROA and uses backward reachability to grow the size of the ROA. Finding a seed ROA can be difficult for general nonlinear systems, but for PWA systems can be easily found due to the dynamics being locally linear. Although reachability-based computation of invariant sets has been explored for PWA dynamical systems before [60], a novelty in our work is pairing these methods with our RPM algorithm so that they can apply to dynamical systems represented as neural networks. Our proposed method for finding ROAs of neural networks follows the backward reachability approach from the literature. In addition, we show that backward reachability may be sped up considerably when restricting RPM to enumerate a connected backward reachable set (a unique feature enabled by the connected walk of the algorithm).
@@ VI-A Control Invariant Sets
In this subsection we discuss a strategy for computing control invariant sets once we have the PWA representation of a dynamical system $\mathbf{x}_{t+1}=F(\mathbf{x}_{t},\mathbf{u}_{t})$. This strategy for computing control invariant sets is implemented in MPT3 specifically for PWA dynamical systems [12]. First, we note that for some state transition functions $F$, there may not exist a control invariant set. Existence depends on the state domain $\mathcal{X}$ and the set of admissible control inputs $\mathcal{U}$, as well as the effect of control inputs on the state transitions given by $F$. If a control invariant set exists, it can be computed via the following recursion which uses backward reachability,
@@ VI-B Regions of Attraction
In this subsection we consider the problem of computing invariant ROAs given a ReLU network that represents an autonomous dynamical system $\mathbf{x}_{t+1}=F(\mathbf{x}_{t})$. One of the most common goals in designing a control policy is to stabilize the system around some equilibrium. By identifying a ROA, we are identifying a set of states which is stabilized.
@@ VI-C An Accelerated Procedure
The backward reachability algorithm as previously presented can be made more efficient if RPM is restricted to only enumerated a connected backward reachable set. Knowing that the fixed point $\mathbf{x}_{eq}$ must be in the ROA, initializing RPM at the fixed point, and then only enumerating a connected backward reachable set of the seed ROA can significantly reduce the computation time. In this case, we can alter the backward reachability algorithm to only add neighbor polyhedra to the working set if they are neighbors of a polyhedron known to be in the backward reachable set.
@@
PWA homeomorphic conditions are given in [71].
@@ VII Examples
We then find a seed ROA for the stable affine system as described in Section VI-B2. Next, we verify the network is a homeomorphism by evaluating the conditions in Section VI-C; from this we know that a connected set in the output space will have a connected preimage. We repeat the network training and these preprocessing steps $10$ times and provide ranges for the number of affine regions and computation time for each step in Table I. We run the longer, backward reachability computation, on only one of the networks (a network with 248 regions).
@@
Number of Affine Regions $93-258$ Regions
Time to Enumerate Affine Regions $1.93-3.05$ Seconds
Time to Check if Homeomorphism $0.04-0.04$ Seconds
Time to Find Fixed Point $0.27-0.33$ Seconds
Time to Find Seed ROA $0.74-2.2$ Seconds
@@
In addition, Table II demonstrates that using RPM with knowledge of the homeomorphism also provides significant performance gains over a state of the art exact reachability algorithm [9]. Although the approach from [9] is clearly faster than the “naive” RPM implementation, this is not the case when using the homeomorphic implementation. Moreover, the state bounds were chosen with prior knowledge of the true ROA, and if we relax the state bounds to $x_{1}\in[-10,10]$, $x_{2}\in[-10,10]$, then the method from [9] enumerates the 15-step ROA in 610 seconds (scaling with the size of the state space), whereas the homeomorphic RPM implementation does not scale with the size of the state space (but rather with the size of the ROA), and thus the solve time remains 157 seconds.
@@
Steps # in ROA # Explored Time (s) # Explored Time-RPM Time-[9]
5 $500$ $594$ $5$ $7362$ $98$ 8
10 $2525$ $2735$ $42$ $33,946$ $627$ 127
15 $6569$ $6898$ $157$ $85,217$ $2280$ 295
20 $12,553$ $12,995$ $413$- -- --
25 $20,949$ $21,554$ $954$- -- --
30 $32,753$ $33,500$ $1956$- -- --
@@ VII-B Pendulum Example
as a discrete-time model where $m=1$kg, $g=9.81$m/s2, $l=1$m, and $\theta$ is measured clockwise from the upright position. We repeat the network training and enumeration of affine regions $10$ times and provide ranges for the number of affine regions and computation time in Table III. We run the longer, invariant set computation, on only one of the networks (a network with 217 regions).
Details of the learned dynamics network are in Table III.
@@
Number of Affine Regions $217-823$ Regions
Time to Enumerate Affine Regions $2.35-4.11$ Seconds
@@ VII-C Verifying Image-Based Control
lIn this example we apply the tools of Section VI to a larger dynamical system that models the closed-loop behavior of a taxiing airplane controlled from camera images. Our goal is to investigate the set of states that are stabilized by a learned image-based controller by modeling the closed-loop dynamical system with a single ReLU network, and applying the accelerated procedure from Section VI.
The specific system we consider was introduced in [13] where a taxiing airplane is regulated to the centerline of a runway via rudder-angle control input based on images from an onboard camera. Because of the image observation, we cannot model the observation from physical principles. Instead, we modify the learned models in [13] to obtain a network that approximates the observed image given a state
@@
Next, we use Algorithm 4 to retrieve the explicit PWA function for $F_{cl}$ over the domain $p\in[-5,5]$ m., $\theta\in[-15^{\circ},15^{\circ}]$. The resulting PWA function has 244,920 regions, took 3 hours to compute, and was not found to be a homeomorphism. Timing for these steps is summarized in Table V. Although the function is not homeomorphic, when applying the accelerated algorithm from Section VI-C, we are still guaranteed that all states within this set asymptotically converge to the stable fixed point.
@@ VIII Conclusions
We proposed the Reachable Polyhedral Marching (RPM) algorithm to incrementally construct an exact PWA representation of a ReLU network. RPM computes an explicit PWA function representation for a given ReLU network that can then be used to quickly find forward and backward reachable sets of ReLU networks. In addition, we demonstrated how, for dynamical systems modeled as ReLU networks, RPM can be leveraged to (i) compute intricate control invariant sets and (ii) certify stability and compute ROAs. RPM enumerates affine regions and reachable sets in an incremental and connected fashion, which we show can lead to massive acceleration when computing ROAs. Lastly, RPM also provides a novel way to determine invertibility of ReLU networks.